OpenStudy (anonymous):

Solve for x: 7 over 8 minus 1 over x equals 3 over 4. x = 1 x = 2 x = 4 x = 8

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 } - \frac{ 1 }{ x } = \frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

Do the follow 7/8-3/4=1/x 7-6/8=1/x 1/8=1/x Now you can try to finish;)

OpenStudy (anonymous):

so we have \[\frac{ 1 }{ 8 } = \frac{ 1 }{ x }\] just to make sure

OpenStudy (anonymous):

x = 1?

OpenStudy (anonymous):

i mean 8

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

thnx i have 2 more can you help with?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

post it here or in a new one?

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

Solve for x: 2 over quantity x plus 2 plus 1 over 5 equals 6 over quantity x plus 5. x = 0 x = –13 x = 13 x = 0 and x = 13

OpenStudy (anonymous):

\[\frac{ 2 }{ x + 2 } + \frac{ 1 }{ 5 } = \frac{ 6 }{ x + 5 }\]

OpenStudy (anonymous):

Use minimum multiple common (5.2)+(x+2)/5x+10= 6/x+5 10+x+2/5x+10=6/x+5 12+x/5x+10=6/x+5 (12+x).(x+5)=6.(5x+10) Now you can try to solve the rest

OpenStudy (anonymous):

im stuck at 62 + x^2 = 25x

OpenStudy (anonymous):

should i subtract x^2 from both sides

OpenStudy (anonymous):

No, now you have to use baskara

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

Look you made a mistake somewhere 12x+60+x^2+5x=30x+60 17x+60+x^2=30x+60 X^2+17x=30x X^2-13x=0 So X(x-13)=0 X=0 Or X=13

OpenStudy (anonymous):

D is right

OpenStudy (anonymous):

ohh okay i see

OpenStudy (anonymous):

can you help me with 1 more>?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

Solve for x: 3 over 3 x plus 1 over quantity x plus 4 equals 10 over 7x . x = –3 x = 3 x = –3 and x = 3 No Solution

OpenStudy (anonymous):

\[\frac{ 3 }{ 3x } + \frac{ 1 }{ x + 4 } = \frac{ 10 }{ 7x }\]

OpenStudy (anonymous):

1/x+4=10/7x-3/3x 1/x+4=10/7x-1/x 1/x+4=10-7/x 1/x+4=3x Try to finish now

OpenStudy (anonymous):

give me 1 sec plz

OpenStudy (anonymous):

1/x+4=3/x Sorry I forgot the slash

OpenStudy (anonymous):

is it -3?

OpenStudy (anonymous):

oh ok let me try again

OpenStudy (anonymous):

idk what the answer is

OpenStudy (anonymous):

X=3x+12 -2x=12 X=-6

OpenStudy (anonymous):

but that not one of my answer choices

OpenStudy (anonymous):

x = –3 x = 3 x = –3 and x = 3 No Solution

OpenStudy (anonymous):

My gosh I've made a mistake 1/x+4=3/7x 7x=3x+12 4x=12 X=3 Sorry about that

OpenStudy (anonymous):

np thnx